OpenStudy (anonymous):

Please help!! Given the function f(x) = 4(2)x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (triciaal):

can you clarify your function?

OpenStudy (triciaal):

rate of change is the slope

OpenStudy (anonymous):

I dont understand what you mean im sorry

OpenStudy (triciaal):

you have f(x) = 4(2)x I don't know what this means. Can you say what it means? I think this should be an exponential curve

OpenStudy (anonymous):

\[f(x)=4(2)^{x}\]

OpenStudy (anonymous):

and yes its supposed to be an exponential curve

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

rate of change section A f(2) - f(1) divided by (2 - 1 ) = rate of change section B = f(4) - f(3) divided by (4 - 3) =

OpenStudy (triciaal):